OpenStudy (anonymous):

if cos0=-40/41 and sin0=9/41 what is the value of tan0 in quadrant 2

OpenStudy (luigi0210):

Let's draw a picture.

OpenStudy (anonymous):

would it be -40/9

OpenStudy (luigi0210):

Not exactly. \[\LARGE tan \theta=\frac{opp}{adj}\]

OpenStudy (***[isuru]***):

use this \[1 + \tan ^{2} \theta = \sec ^{2}\theta \] it gives u the most advance answer

OpenStudy (haseeb96):